# FlashMD: long-stride, universal prediction of molecular dynamics

### v1

@@ Appendix H Liquid argon benchmark
Appendix I Water simulations based on the q-TIP4P/f model
@@
Appendix J Simulation details
@@
simulations at 450 K and 1 bar were performed with a single dipeptide solvated by 200 molecules of water in a cubic cell of Å3, as done in Morrone et al. [76]. The initial configuration of the system was randomly initialized with packmol [97], and the starting conformations of the dipeptide were taken from Ref. [98]. Simulations were performed for PET-MAD MLIP at 0.5 fs strides, universal FlashMD with 8 fs strides, and universal FlashMD with 16 fs strides. Langevin thermostat was coupled to the system with fs, and an isotropic Bussi-Zykova-Parrinello (BZP) barostat [83] was used with fs, including a Langevin thermostat coupled to the cell parameters with fs. For each MD engine, 10 simulations were parallelly performed with different starting configurations to optimize sampling. Despite the difference in time strides, equivalent number of snapshots were sampled across the simulation setups to ensure a fair comparison of the resulting free energy surfaces.
@@
Appendix K Computational resources

### v2

@@ Appendix H Liquid argon benchmark (MDNet)
Appendix I SPC/E water benchmark (TrajCast)
TrajCast [63] published three datasets for the direct learning of molecular dynamics. Here, we compare our accuracies with the accuracies reported in Thiemann et al. [63], using the SPC/E water dataset they provide.
| Architecture | TrajCast | FlashMD |
| Displacements | 0.17 | 0.17 |
| Momenta | 0.37 | 0.22 |
The FlashMD results here make use of a slightly improved architecture compared to the one used to generate the results shown in this work. The same architecture was used to train the r2SCAN FlashMD models we currently recommend, and it corresponds to the FlashMD implementation in metatrain [82]. The architecture used to produce the results in this work, for example, yields a momentum MAE of 0.52%, indicating that FlashMD and TrajCast have similar accuracies and that relatively minor tweaks can tip the numbers in favor of one or the other. In general, we always found our models to show clear signs of underfitting, showing that larger models and/or longer training times are generally beneficial when training models for the direct prediction of molecular dynamics trajectories, especially when compared to machine-learned interatomic potentials. The design of more accurate models will be a crucial challenge to achieve near-quantitative results using direct models for molecular dynamics in the future.
Appendix J Water simulations based on the q-TIP4P/f model
@@ Appendix I SPC/E water benchmark (TrajCast)
Appendix K Simulation details
@@
simulations at 450 K and 1 bar were performed with a single dipeptide solvated by 200 molecules of water in a cubic cell of Å3, as done in Morrone et al. [79]. The initial configuration of the system was randomly initialized with packmol [105], and the starting conformations of the dipeptide were taken from Ref. [106]. Simulations were performed for PET-MAD MLIP at 0.5 fs strides, universal FlashMD with 8 fs strides, and universal FlashMD with 16 fs strides, for a total duration of 1 ns. Langevin thermostat was coupled to the system with fs, and an isotropic Bussi-Zykova-Parrinello (BZP) barostat [92] was used with fs, including a Langevin thermostat coupled to the cell parameters with fs. For each MD engine, 10 simulations were parallelly performed with different starting configurations to optimize sampling. Despite the difference in time strides, equivalent number of snapshots were sampled across the simulation setups to ensure a fair comparison of the resulting free energy surfaces.
@@
Appendix L Computational resources